Mike's age is 3/4 of Mary's age. In 5 yrs, Mike's age will be 7/9 Mary's age. How old is Mike now?
Answer by alicealc(293)

Mike's age = A
Mary's age = B
A = 3/4 * B
In 5 years:
A + 5 = 7/9 * (B + 5)
3/4 * B + 5 = 7/9 * B + 35/9
3/4 * B - 7/9 * B = 35/9 - 5
(27/36 - 28/36) * B = 35/9 - 45/9
-1/36 * B = -10/9
B = -10/9 * -36/1
B = 360/9 = 40
A = 3/4 * 40 = 30
So, Mike's age now is 30 years old